Two angles form a linear pair. One angle is 4 less than three times the other angle. What is the measure of the obtuse angle?
Answer by cleomenius(959)

3x - 4 + x = 180
4x = 184
x = 46 the first angle.
3(46) = 138 - 4 = 134 degrees.
134 + 46 = 180 degrees.
This does check.
Cleomenius.